OpenStudy (anonymous):

the sum of three consecutive integers is 294. Find the integers.

OpenStudy (texaschic101):

x = 1st integer x + 1 = 2nd integer x + 2 = 3rd integer (x ) + (x + 1) + (x + 2) = 294 -- combine like terms 3x + 3 = 294 -- subtract 3 from both sides 3x = 294 - 3 3x = 291 -- divide both sides by 3 x = 97 x + 1 = 97 + 1 = 98 x + 2 = 97 + 2 = 99 your numbers are : 97,98,99

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

I think I did something wrong because I got 96,98, 100

OpenStudy (anonymous):

Nevermind, I see what I did incorrectly. I solved the problem for EVEN numbers.

OpenStudy (anonymous):

Thank you, man.